# Fundamentals and Advances in Transverse Thermoelectrics

### v1

@@ VI.4 Thermoelectric permanent magnet
  • Zero-magnetic-field operation of the anomalous Nernst and Ettingshausen effects [60, 61, 160],
@@
Permanent magnets are not homogeneous materials but rather composite materials or engineered ensembles of interfaces with micrometer- to sub-micrometer-scale structures [Fig. 11(b)]. The goal of conventional permanent magnet development has been to achieve large maximum energy product $(BH)_{\rm max}$, and the optimization of microstructures has significantly improved the coercive force, remanent magnetization, and their temperature-dependent characteristics [162]. The design guidelines for the thermoelectric permanent magnets are completely different from those for conventional magnets; it is necessary to design and control microstructures to improve the transverse thermoelectric conversion characteristics. For example, by incorporating thermoelectric semiconductors into the microstructure of permanent magnets, the appearance of the Seebeck-driven ordinary/anomalous Hall effects in the magnets is expected. Introducing spin-orbit interaction near the interface between ferromagnetic phases and nonmagnetic grain-boundary phases may result in the superposition of the transverse thermopower due to spin-Seebeck/spin-dependent-Seebeck-driven ISHEs [see the right magnified illustration in Fig. 11(b)]. Furthermore, by appropriately designing the grain size and distribution, it may be possible to simultaneously reduce the phonon thermal conductivity [43, 163] [see the left illustration in Fig. 11(b)]. In other words, by optimally designing the microstructures of the thermoelectric permanent magnets, it is possible to independently engineer the transverse thermopower, electrical conductivity, and thermal conductivity, thereby significantly improving the figure of merit and energy conversion efficiency. To achieve this, it will be necessary to go beyond mere material exploration, and instead focus on the design of hierarchical structures and the control of phase boundaries within permanent magnets, with the aid of microstructure analysis techniques and precise transport property measurements. Microstructure engineering of magnetic materials aimed at realizing high-performance transverse thermoelectric conversion has already begun [43, 56, 58, 59, 164].
@@ Acknowledgments
The authors thank Y. Goto, J. P. Heremans, K. Hirata, Y. Lee, Y. Mizuguchi, S. Mori, M. Murata, H. Nagano, Y. Nakanishi, K. Oyanagi, S. J. Park, Y. Sakuraba, Y. Sato, H. Sepehri-Amin, A. Takahagi, T. Yagi, and W. Zhou for valuable discussions and many collaborators and group members for preliminary reviews of the manuscript. The previously reported experimental results introduced in this article were obtained from the collaboration with H. Masuda, T. Ohkubo, Y. Sakuraba, T. Seki, H. Sepehri-Amin, K. Takanashi, T. Yagi, and W. Zhou. This work was mainly supported by ERATO “Magnetic Thermal Management Materials” (grant no. JPMJER2201) from Japan Science and Technology Agency.

### v2

@@ VI.4 Thermoelectric permanent magnet
  • Zero-magnetic-field operation of the anomalous Nernst and Ettingshausen effects [63, 64, 171, 109],
@@
Permanent magnets are not homogeneous materials but rather composite materials or engineered ensembles of interfaces with micrometer- to sub-micrometer-scale structures [Fig. 11(b)]. The goal of conventional permanent magnet development has been to achieve large maximum energy product $(BH)_{\rm max}$, and the optimization of microstructures has significantly improved the coercive force, remanent magnetization, and their temperature-dependent characteristics [173]. The design guidelines for the thermoelectric permanent magnets are completely different from those for conventional magnets; it is necessary to design and control microstructures to improve the transverse thermoelectric conversion characteristics. For example, by incorporating thermoelectric semiconductors into the microstructure of permanent magnets, the appearance of the Seebeck-effect-driven ordinary/anomalous Hall effects in the magnets is expected. Introducing spin-orbit interaction near the interface between ferromagnetic phases and nonmagnetic grain-boundary phases may result in the superposition of the transverse thermopower due to SSE/SdSE-driven ISHEs [see the right magnified illustration in Fig. 11(b)]. Furthermore, by appropriately designing the grain size and distribution, it may be possible to simultaneously reduce the phonon thermal conductivity [46, 174] [see the left illustration in Fig. 11(b)]. In other words, by optimally designing the microstructures of the thermoelectric permanent magnets, it is possible to independently engineer the transverse thermopower, electrical conductivity, and thermal conductivity, thereby significantly improving the figure of merit and energy conversion efficiency. To achieve this, it will be necessary to go beyond mere material exploration, and instead focus on the design of hierarchical structures and the control of phase boundaries within permanent magnets, with the aid of microstructure analysis techniques and precise transport property measurements. Microstructure engineering of magnetic materials aimed at realizing high-performance transverse thermoelectric conversion has already begun [46, 59, 61, 62, 175].
@@ Acknowledgments
The authors thank Y. Goto, J. P. Heremans, K. Hirata, M. Hirschberger, Y. Lee, Y. Mizuguchi, S. Mori, M. Murata, H. Nagano, Y. Nakanishi, K. Oyanagi, S. J. Park, Y. Sakuraba, Y. Sato, H. Sepehri-Amin, A. Takahagi, T. Yagi, and W. Zhou for valuable discussions and many collaborators and group members for preliminary reviews of the manuscript. The previously reported experimental results introduced in this article were obtained from the collaboration with H. Masuda, T. Ohkubo, Y. Sakuraba, T. Seki, H. Sepehri-Amin, K. Takanashi, T. Yagi, and W. Zhou. This work was mainly supported by ERATO “Magnetic Thermal Management Materials” (grant no. JPMJER2201) from Japan Science and Technology Agency.